If five times the smaller of two consecutive integers is
added to three times the larger, the result is 59. Find the integers.

Answer by checkley77(12844)

Let x & x+1 be the 2 integers.
5x+3(x+1)=59
5x+3x+3=59
8x=59-3
8x=56
x=56/8
x=7 ans.
x+1=8 ans.
Proof:
5*7+3*88=57
35+24=59
59=59